OpenStudy (anonymous):

solve the equation 2+ 4 cos^2 2x^(degree) giving your interval from -180 to 180

OpenStudy (sirm3d):

\[\large 2 + 4 \cos ^2 2x = ? \]where \[\large -180\text O<x \le 180^\text O\]

OpenStudy (anonymous):

2x where x has a degree on top